There is a disk rotating with a constant angular speed of 18 rad/sec. At some instant the disk begins to experience a constant friction resulting in an angular acceleration of 1.4 rad/sec2

Find:

1) the time (in sec) it takes the disk to come to a complete stop due to friction

2) total number of complete revolutions the disk undergoes before coming to a complete stop due to friction;

3) the time (in sec) it takes the disk to complete 8 full revolutions (be careful here!).
Answer #1

1)

w = wo + \alpha t

0 = 18 + ( - 1.4) t

t = 18 / 1.4

t = 12.86 seconds

______________

2)

w2 - wo2 = 2\alpha\theta

02 - 182 = 2 * -1.4 * \theta

\theta = 115.7 radians

so,

Number of turns = \theta / 2* pi

N = 18.416 ( or 18 full revolutions)

________________

3)

8 full revolutions means 8 * 2 * pi = 50.265 rad

so,

\theta = wot - 1/2\alphat2

50.265 = 18t - 0.7t2

0.7t2 - 18t + 50.265 = 0

t = 18 +/ - 13.54 / 2 * 0.7

t = 3.186 seconds